# BFS-Prover: Scalable Best-First Tree Search for LLM-based Automatic Theorem Proving

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have spurred growing interest in automatic theorem proving using Lean4, where effective tree search methods are crucial for navigating proof search spaces. While the existing approaches primarily rely on value functions and Monte Carlo Tree Search (MCTS), the potential of simpler methods like Best-First Search (BFS) remains underexplored. This paper investigates whether BFS can achieve competitive performance in large-scale theorem proving tasks. We present BFS-Prover, a scalable expert iteration framework, featuring three key innovations. First, we implement strategic data filtering at each expert iteration round, excluding problems solvable via beam search node expansion to focus on harder cases. Second, we improve the sample efficiency of BFS through Direct Preference Optimization (DPO) applied to state-tactic pairs automatically annotated with compiler error feedback, refining the LLM’s policy to prioritize productive expansions. Third, we employ length normalization in BFS to encourage exploration of deeper proof paths. BFS-Prover achieves a score of $71.31$ on the MiniF2F test set and therefore challenges the perceived necessity of complex tree search methods, demonstrating that BFS can achieve competitive performance when properly scaled.
@@ 1 Introduction
Automatic theorem proving (ATP) in formal languages has recently become a critical benchmark for evaluating the reasoning capabilities of large language models (LLMs). By encoding mathematical problems into formal systems like Lean4, ATP enables the generation of machine-verified proofs for complex mathematical propositions, ensuring logical correctness [10, 14]. Despite the remarkable success of LLMs in natural language mathematics and reasoning tasks [7, 11], theorem proving in formal languages presents unique challenges [24]. Unlike informal reasoning, formal systems require strict adherence to syntax and semantics, as well as the ability to generate valid steps within a highly constrained formal framework. Additionally, the action (tactic) space in ATP is vast, as each proof state can lead to numerous potential tactics, making the search process for valid proofs computationally intensive [13].
@@
Best-First Search (BFS) [12] offers a simpler and more lightweight alternative to MCTS by prioritizing expansions based on accumulated log probabilities from the current node to the root. While its straightforward nature and computational simplicity make it attractive, BFS is often considered suboptimal for theorem proving in the existing literature [20, 9, 22] due to the following hypothesis:
  • Lack of exploration: BFS prioritizes high-probability paths, overlooking less likely but valid solutions. Without exploration mechanisms like upper confidence bounds or value functions, it struggles to balance exploiting promising nodes and exploring diverse paths.
@@ 1.1 Our Contributions
This paper challenges the prevailing perception that BFS is inherently unsuitable for large-scale ATP. We present BFS-Prover system, which transforms BFS into a simple yet powerful algorithm through targeted scaling strategies. Our key contributions are as follows.
@@
  • Empirical Validation on MiniF2F. BFS-Prover achieves a score of 70.08 on the MiniF2F test set, demonstrating that BFS can scale to competitive levels in ATP. This result highlights the viability of BFS as a lightweight yet powerful alternative to more complex methods like MCTS and value functions.
@@ 2 The BFS-Prover system
Figure 1: An overview of the BFS-Prover system. The figure illustrates two key components: a) Best-First Tree Search (BFS) mechanism for automatic theorem proving, showing how the system explores proof paths by expanding nodes based on accumulated probabilities, with different node types (proof finish, error, intermediate, and unexplored) represented via color coding. The identified proof path demonstrates a successful state-tactic sequence. b) The training data accumulation process, depicting how the system leverages both successful proofs for Supervised Fine-Tuning (SFT) and compiler error feedback for Direct Preference Optimization (DPO). The system integrates LeanDojo as a python interface for Lean4 and employs a language model as the policy network for tactic generation.
@@ 2.3 Expert Iteration
    Direct Preference Optimization (DPO): Instead of re-training the policy LLM using SFT, DPO can be applied as an alternative method to refine the current policy model by leveraging the on-policy Lean error data collected in step 2. Along the proof path, certain generated tactics, expanded from a proof state but not part of the verified proof path, result in Lean compiler errors. These erroneous tactics naturally serve as negative examples when compared to their corresponding valid proof tactics, forming preference tuples. Specifically, for a state $s$ on a proof path, we may form pairs $(a_{w},a_{l})$ where $a_{w}$ is the tactic on the proof path and $a_{l}$ is an error-causing tactic expanded from $s$ if exists; see Fig. 1 for an illustration. The DPO loss is then defined as:
    $\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{(s,a_{w},a_{l})}\left[\log\sigma(\beta(r_{\theta}(s,a_{w})-r_{\theta}(s,a_{l})))\right],$
    where $r_{\theta}(s,a)=\log(p_{\theta}(a|s))-\log(p_{\text{ref}}(a|s))$ is the log probability ratio between the policy model $p_{\theta}$ and reference model $p_{\text{ref}}$, $\beta$ is a KL regularization parameter controlling the sharpness of the learned preferences, and $\sigma$ is the sigmoid function. By explicitly incorporating these negative signals that are absent in SFT, DPO sharpens the LLM’s output distribution, thereby improving the sample efficiency and scalability of BFS within the expert iteration framework.
@@ 3.1 Model, Data, and Training Setup
Base model and initial training data. For illustration purposes, we use Qwen2.5-Math-7B [23] as the base model for fine-tuning the policy LLM in BFS-Prover. In order to initialize the expert iteration process, we leverage proof data extracted by LeanDojo [25] from Mathlib [10], which serves as the cold-start dataset. As expert iteration progresses, we further incorporate datasets from Lean-Github [21], a compilation of Lean4 repositories from GitHub, and Lean-Workbook [26], which focuses on Olympia-level algebra and analysis. These datasets span a diverse range of mathematical topics and formal reasoning tasks, equipping the policy model with the foundational skills necessary for effective tactic generation and proof navigation.
@@
Training setup. At each expert iteration round, we obtain a new policy LLM through either SFT or DPO. For SFT, we conduct training over 3 epochs using cosine learning rate decay, initializing at $2\times 10^{-5}$ and decaying to $1\times 10^{-6}$, with a global batch size of $128$. For DPO, we perform a single epoch of training with cosine learning rate decay, beginning at $5\times 10^{-6}$ and decaying to $5\times 10^{-7}$, utilizing a global batch size of $16$, while maintaining a KL regularization parameter of $\beta=10$. All training is performed on A100 80GB GPUs. The selection between SFT and DPO depends on the volume of data generated in each iteration: SFT is implemented when an iteration yields substantial new data, whereas DPO is preferred for iterations with limited data generation, leveraging its sample efficiency through the incorporation of negative examples.
BFS configuration in expert iteration. We set the length normalization parameter $\alpha$ to $0.0$ throughout. During the beam search filtering stage, we utilize a beam width of $32$ to identify easily solvable theorems. In the subsequent data collection phase, we employ temperature-based sampling with a temperature of $1.0$, nucleus sampling parameter $1.0$, and a sampling width of either $4$ or $8$ to explore diverse proof paths.
@@ 3.2 Distributed Best-First Search Infrastructure
To enable efficient large-scale proof search, we implement a distributed system using Ray for parallel theorem proving across multiple machines, each equipped with $8$ A100 80GB GPUs and $128$ CPU cores. The target theorems are evenly divided across machines, with each machine running an independent proving pipeline. The system is composed of three main components:
  • GPU-Based Policy LLM Pool: Each local machine deploys 8 instances of our 7B policy LLM, each powered by an asynchronous vLLM engine running on a dedicated A100 GPU. These instances form a shared pool that processes concurrent tactic generation requests.
  • CPU-Based Prover Pool: Each local machine runs $96$ concurrent prover instances, reserving remaining CPU cores for system operations. Each prover instance executes independent best-first search runs for its assigned theorems. In order to achieve balanced GPU utilization, provers distribute their requests across the policy LLM instances in a round-robin fashion based on their index modulo $8$. Each prover asynchronously interacts with its assigned policy LLMs and LeanDojo environments.
  • Asynchronous Orchestration: The overall system leverages asyncio to manage the high-concurrency workflow between provers and policy LLMs. Both the policy LLM and prover pools are implemented as Ray actors, enabling dynamic resource management through Ray’s runtime system. To ensure system responsiveness, we enforce timeout thresholds for both tactic execution (via LeanDojo) and model inference (via vLLM).
This design achieves near-linear scaling through efficient theorem distribution across machines while maximizing hardware utilization within each machine without cross-machine communication overhead.
@@ 3.3 Distribution Shift in Expert Iteration
Proof level. A critical indicator of the effectiveness of BFS-Prover is its ability to discover deep proofs. We define proof length as the number of tactics found by our system to complete a proof. We observe that the distribution of proof length at each expert iteration round often exhibits a Gaussian or mixture of Gaussians, reflecting the diversity of theorem complexities within the formal statement corpus. Intriguingly, as expert iteration progresses, the mean proof length tends to increase, indicating that BFS can find increasingly deeper and more challenging proofs as the policy LLM improves; see Fig. 2 for an illustration. This observation highlights the effectiveness of the expert iteration framework and the scalability of BFS.
Tactic level. Beyond proof-level evolution, we also observe interesting distributional shifts at the tactic level throughout expert iteration; see Fig. 3. Notably, the BFS-Prover system maintains a diverse range of tactic lengths without collapsing to a narrow distribution - a common failure mode in reinforcement learning where models tend to converge to a limited set of high-reward actions [19]. Instead, we observe a moderate but meaningful shift from very simple tactics (1-10 tokens) towards more commonly used tactical patterns (11-50 tokens). This shift suggests that through expert iteration, BFS-Prover learns to generate more sophisticated tactics while maintaining the capability to use simpler ones when appropriate. The preservation of tactic diversity is crucial for effective theorem proving, as different proof states require different levels of tactical complexity, from simple term rewriting to complex algebraic manipulations.
Figure 3: Sample evolution of tactic token length distribution across expert iteration rounds. The plot shows the log-scaled percentage of tactics falling into different token length bins. We have two observations: (1) robustness against mode collapse during expert iteration, evidenced by the maintained diversity across token lengths, and (2) a gradual shift from an emphasis on very short tactics (111110101010 tokens, decreasing by about $2.3\%$) to a higher prevalence of moderate-length tactics (1111111150505050 tokens, with incremental increases ranging from approximately $1.27\%$ to $0.06\%$). This evolution reflects the improved ability of BFS-Prover to generate more complex, informative tactics while still retaining the capacity to produce concise ones when needed.
@@ 3.4 Results on MiniF2F
In this section, we present the performance of BFS-Prover on the MiniF2F test benchmark [27], which is a widely recognized dataset for evaluating formal mathematics systems. MiniF2F consists of a diverse collection of formalized high-school and competition-level mathematics problems. The policy LLM checkpoint for tactic generation used for evaluation was obtained by performing SFT on all state-tactic pairs accumulated through the expert iteration pipeline of BFS-Prover, followed by an additional round of DPO refinement using Lean error signals as described in Section 2.3.
@@ 3.4.1 Comparison with the State of the Art
Fixed tactic generation budget. We compare BFS-Prover developed in this work against leading theorem provers in the literature, including DeepSeek-Prover-V1.5 [22], InternLM2.5-StepProver [20], and HunyuanProver [9]. For fair comparison, we first evaluate under a fixed tactic generation budget, defined as $K\times W\times N$, where $K$ is the number of passes, $W$ is the expansion width, and $N$ is the maximum number of proof state expansions per pass. During BFS proof search, BFS-Prover uses a sampling temperature of $1.1$, expansion width of $2$, and length normalization factor of $\alpha=0.7$ defined in (1) to maximize the diversity of explored proof paths across passes. As shown in Table 1, BFS-Prover achieves state-of-the-art performance without requiring either a critic model (value function) or MCTS-based search. We note that the incorporation of a critic model effectively doubles the number of inference calls since each state expansion requires both policy and value computations.
[…]
Prover System Policy Model Critic Model Tree Search Tactic Budget MiniF2F-test
@@
DeepSeek-Prover-V1.5 [22] 7B No MCTS $32\times 16\times 400$ 63.5%
@@
InternLM2.5-StepProver [20] 7B Yes BFS $256\times 32\times 600$ 65.9%
HunyuanProver [9] 7B Yes BFS $600\times 8\times 400$ 68.4%
BFS-Prover (this work) 7B No BFS $2048\times 2\times 600$ 68.85%
BFS-Prover (this work) 7B No BFS accumulative 71.31%
@@
3.4.2 Scaling Law of BFS and Advantage of DPO
In this subsection, we study how the performance of BFS-Prover scales with the number of proof search passes, as well as the role of DPO in enhancing the system. We perform a total of pass@2048 and evaluate intermediate performance at pass@64, pass@128, pass@256, and pass@1024, where the confidence band for each intermediate pass count is computed by sampling multiple pass@64 runs. The experimental results are presented in Fig. 4, where $x$-axis is in the log scale. In the following, we observe several key characteristics of BFS scaling in formal theorem proving.
  • Both SFT and SFT+DPO training methods demonstrate logarithmic scaling with increasing passes. Specifically, SFT improves from 62.60% to 67.21% and SFT+DPO from 63.23% to 68.85% when scaling passes from 64 to 2048, showing consistent but diminishing returns with doubled computational budget.
  • The DPO-enhanced model consistently achieves a performance advantage over the SFT baseline, with the gap growing from 0.63% at 64 passes to 1.64% at 2048 passes. This widening gap suggests that the benefits of DPO become more pronounced with increased computational budget.
  • Examining the min-max ranges, both methods exhibit similar variance in performance (around 2-3% range). This suggests that while DPO improves overall performance, it maintains comparable stability in proof search to the SFT baseline.
@@ 3.4.2 Scaling Law of BFS and Advantage of DPO
Figure 4: Evolution of proof accuracies with increasing number of proof search passes on the MiniF2F test. We compare two training methods: (blue) only using SFT on all state-tactic pairs accumulated in the expert iteration pipeline of BFS-Prover; and (red), on top of (blue), perform another round of DPO refinement using negative tactic examples described in Section 2.3. It can be seen that DPO achieves significant improvement in terms of both sample efficiency and overall accuracy.
4 Conclusion

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have spurred growing interest in automatic theorem proving using Lean4, where effective tree search methods are crucial for navigating the underlying large proof search spaces. While the existing approaches primarily rely on value functions and/or Monte Carlo Tree Search (MCTS), the potential of simpler methods like Best-First Tree Search (BFS) remains underexplored. In this paper, we investigate whether BFS can achieve competitive performance in large-scale theorem proving tasks. We present BFS-Prover, a scalable expert iteration framework, featuring three key innovations. First, we implement strategic data filtering at each expert iteration round, excluding problems solvable via beam search node expansion to focus on harder cases. Second, we improve the sample efficiency of BFS through Direct Preference Optimization (DPO) applied to state-tactic pairs automatically annotated with compiler error feedback, refining the LLM’s policy to prioritize productive expansions. Third, we employ length normalization in BFS to encourage exploration of deeper proof paths. BFS-Prover achieves a state-of-the-art score of $72.95\%$ on the MiniF2F test set and therefore challenges the perceived necessity of complex tree search methods, demonstrating that BFS can achieve competitive performance when properly scaled. To facilitate further research and development in this area, we have open-sourced our model at https://huggingface.co/bytedance-research/BFS-Prover.
@@ 1 Introduction
Automatic theorem proving (ATP) in formal languages has recently become a critical benchmark for evaluating the reasoning capabilities of large language models (LLMs). By encoding mathematical problems into formal systems like Lean4, ATP enables the generation of machine-verified proofs for complex mathematical propositions, ensuring logical correctness [12, 16]. Despite the remarkable success of LLMs in natural language mathematics and reasoning tasks [9, 13], theorem proving in formal languages presents unique challenges [28, 23, 6, 22]. Unlike informal reasoning, formal systems require strict adherence to syntax and semantics, as well as the ability to generate valid steps within a highly constrained formal framework. Additionally, the action (tactic) space in ATP is vast, as each proof state can lead to numerous potential tactics, making the search process for valid proofs computationally intensive [15].
@@
Best-First Tree Search (BFS) [14] offers a simpler and more lightweight alternative to MCTS by prioritizing expansions based on accumulated log probabilities from the current node to the root. While its straightforward nature and computational simplicity make it attractive, BFS is often considered suboptimal for theorem proving in the existing literature [24, 11, 26] due to the following hypothesis:
  • Lack of efficient exploration: BFS prioritizes high-probability paths, overlooking less likely but valid solutions. Without exploration mechanisms like upper confidence bounds or value functions, it struggles to balance exploiting promising nodes and exploring diverse paths.
@@ 1.1 Our Contributions
This paper challenges the prevailing perception that BFS is inherently unsuitable for large-scale ATP. We present the BFS-Prover system, which transforms BFS into a simple yet powerful algorithm through targeted scaling strategies. Our key contributions are as follows.
@@
  • Empirical results on MiniF2F. BFS-Prover achieves an accumulative score of 72.95% on the MiniF2F test set, surpassing all state-of-the-art theorem provers in the literature, including DeepSeek-Prover-V1.5 [26], InternLM2.5-StepProver [24], and HunyuanProver [11]. This result underscores that BFS-Prover can scale to competitive levels in ATP while maintaining a lightweight design that avoids the complexities of MCTS and value functions.
@@ 2 The BFS-Prover system
Figure 1: An overview of BFS-Prover. The figure illustrates two key components: a) Best-First Tree Search (BFS) mechanism for automatic theorem proving, showing how the system explores proof paths by expanding nodes based on accumulated probabilities, with different node types (proof finish, error, intermediate, and unexplored) represented via color coding. The identified proof path demonstrates a successful state-tactic sequence. b) The training data accumulation process, depicting how the system leverages both successful proofs for Supervised Fine-Tuning (SFT) and compiler error feedback for Direct Preference Optimization (DPO). The system integrates LeanDojo as a python interface for Lean4 and employs a language model as the policy network for tactic generation.
@@ 2.3 Expert Iteration
    Direct Preference Optimization (DPO): Instead of re-training the policy LLM using SFT, DPO can be applied as an alternative method to refine the current policy model by leveraging the on-policy Lean error data collected in step 2. Along the proof path, certain generated tactics, expanded from a proof state but not part of the verified proof path, result in Lean compiler errors. These erroneous tactics naturally serve as negative examples when compared to their corresponding valid proof tactics, forming preference tuples. Specifically, for a state $s$ on a proof path, we may form pairs $(a_{w},a_{l})$ where $a_{w}$ is the tactic on the proof path and $a_{l}$ is an error-causing tactic expanded from $s$ if exists; see Fig. 1 for an illustration. The DPO loss $\mathcal{L}_{\text{DPO}}(\theta)$ [17] is then defined as
    $-\mathbb{E}_{(s,a_{w},a_{l})}\!\left[\log\sigma(\beta(r_{\theta}(s,a_{w})-r_{\theta}(s,a_{l})))\right]\!,$
    where the differemce, i.e., implicit reward, $r_{\theta}(s,a)=\log(p_{\theta}(a|s))-\log(p_{\text{ref}}(a|s))$ is the log probability ratio between the policy model $p_{\theta}$ and reference model $p_{\text{ref}}$, $\beta$ is a KL regularization parameter controlling the sharpness of the learned preferences, and $\sigma$ is the sigmoid function. By explicitly incorporating these negative signals that are absent in SFT, DPO sharpens the LLM’s output distribution, thereby improving the sample efficiency and scalability of BFS within the expert iteration framework.
@@ 3.1 Model, Data, and Training Setup
Base model and initial training data. For illustration purposes, we use Qwen2.5-Math-7B [27] as the base model for fine-tuning the policy LLM in BFS-Prover. In order to properly initialize the expert iteration process, we leverage proof data extracted by LeanDojo [29] from Mathlib [12], which serves as the cold-start dataset. As expert iteration progresses, we further incorporate state-tactic datasets from Lean-Github [25], a compilation of Lean4 repositories from GitHub, and Lean-Workbook [30], which focuses on Olympia-level algebra and analysis. These datasets span a diverse range of formal mathematics topics, equipping the policy LLM with the foundational skills necessary for effective tactic generation and proof navigation.
@@
Training setup and procedures. At each expert iteration round, we obtain a new policy LLM through either SFT or DPO based on the procedures described in Section 2.3. For SFT, we conduct training over 3 epochs using cosine learning rate decay, initializing at $2\times 10^{-5}$ and decaying to $1\times 10^{-6}$, with a global batch size of $128$. For DPO, we perform a single epoch of training with cosine learning rate decay, beginning at $5\times 10^{-6}$ and decaying to $5\times 10^{-7}$, utilizing a global batch size of $16$, while maintaining a KL regularization parameter of $\beta=10$. All training is performed on A100 80GB GPUs. The selection between SFT and DPO at each expert iteration round depends on the volume of data generated in each iteration: SFT is implemented when substantial new data has been generated during the past iteration(s), whereas DPO is preferred for iterations with limited data generation, leveraging its sample efficiency through the incorporation of negative examples.
BFS configuration in expert iteration. We set the length normalization parameter $\alpha$ to $0.0$ throughout, to minimize inductive bias in expert iteration. During the beam search filtering stage, we utilize a beam width of $32$ to identify easily solvable theorems. In the subsequent data collection phase, we employ temperature-based sampling with a temperature of $1.0$, nucleus sampling parameter $1.0$, and a sampling width of $2$, $4$, or $8$ to explore diverse proof paths.
@@ 3.2 Distributed Best-First Search Infrastructure
To enable efficient large-scale parallel proof search, we implement a distributed system using Ray for distributed theorem proving across multiple machines, each equipped with $8$ A100 80GB GPUs and $128$ CPU cores. The target theorems are evenly divided across machines, with each machine running an independent proving pipeline. The system is composed of three main components:
  • GPU-Based Policy LLM Pool: Each local machine deploys 8 instances of our 7B policy LLM, each powered by an asynchronous vLLM engine [7] running on a dedicated A100 GPU. These instances form a shared pool that processes concurrent tactic generation requests.
  • CPU-Based Prover Pool: Each machine runs $96$ concurrent prover instances, reserving remaining CPU cores for standard system operations. Each prover instance executes independent BFS runs for its assigned theorems. In order to achieve balanced GPU utilization, provers distribute their requests across the policy LLM instances in a round-robin fashion based on their index modulo $8$. Each prover asynchronously interacts with its assigned policy LLMs and LeanDojo environments.
  • Asynchronous interactions: The overall distributed search system leverages asyncio to manage the high-concurrency workflow between provers and policy LLMs. Both the policy LLM and prover pools are implemented as Ray actors, enabling dynamic resource management through Ray’s runtime system. To ensure system responsiveness, we enforce timeout thresholds for both tactic execution (via LeanDojo) and model inference (via vLLM).
This distributed infrastructure design achieves near-linear scaling through efficient theorem distribution across machines while maximizing hardware utilization within each machine without cross-machine communication overhead.
@@
Figure 3: Sample evolution of tactic token length distribution across expert iteration rounds. The plot shows the log-scaled percentage of tactics falling into different token length bins. We have two observations: (1) robustness against mode collapse during expert iteration, evidenced by the maintained diversity across token lengths, and (2) a gradual but meaningful shift from an emphasis on very short tactics (111110101010 tokens, decreasing by about $2.3\%$) to a higher prevalence of moderate-length tactics (1111111150505050 tokens, with incremental increases ranging from approximately $1.27\%$ to $0.06\%$). This evolution reflects the improved ability of BFS-Prover to generate more complex and informative tactics while still retaining the capacity to produce concise ones when needed.
@@ 3.3 Distribution Shift in Expert Iteration
In this subsection, we discuss and illustrate how distribution shifts emerge at both the proof level and the tactic level throughout expert iteration, shedding light on the progressively improving capabilities of BFS-Prover in theorem proving.
Proof level. A critical indicator of the effectiveness of a prover system, such as BFS-Prover, is its ability to discover deep proofs. We define proof length as the number of tactics found by a system to complete a proof. We observe that the distribution of proof length at each expert iteration round often exhibits a Gaussian or mixture of Gaussians, reflecting the diversity of theorem complexities within the formal statement corpus. Intriguingly, as expert iteration progresses, the mean proof length tends to increase, indicating that BFS can find increasingly deeper and more challenging proofs as the policy LLM improves; see Fig. 2 for an illustration. This observation highlights the effectiveness of the expert iteration framework and the scalability of BFS, i.e., its ability to progressively tackle more complex proofs by leveraging iterative policy refinement and improved search capabilities.
Tactic level. Beyond proof-level evolution, we also observe interesting distributional shifts at the tactic level throughout expert iteration; see Fig. 3. Notably, the policy LLM in the BFS-Prover system maintains a diverse range of tactic lengths without collapsing to a narrow distribution - a common failure mode in reinforcement learning where models tend to converge to a limited set of high-reward actions [21]. Instead, we observe a moderate but meaningful shift from very simple tactics (1-10 tokens) towards more commonly used tactical patterns (11-50 tokens). This shift suggests that through expert iteration, the policy LLM in BFS-Prover learns to generate more sophisticated tactics while maintaining the capability to use simpler ones when appropriate. The preservation of tactic diversity is crucial for effective theorem proving, as different proof states require different levels of tactical complexity, from simple term rewriting to complex algebraic manipulations.
@@ 3.4 Results on MiniF2F
In this subsection, we discuss BFS-Prover’s performance on the MiniF2F test benchmark [31] that is a widely recognized dataset for evaluating formal mathematics systems. MiniF2F is consisted of a diverse collection of formalized competition-level mathematics problems. The policy LLM checkpoint for tactic generation used for evaluation was obtained by performing SFT on all state-tactic pairs accumulated through $10$ rounds of expert iteration pipeline of BFS-Prover, followed by an additional round of DPO refinement using Lean error signals as described in Section 2.3.
@@ 3.4.1 Comparison with the State of the Art
We now compare BFS-Prover developed in this work against leading theorem provers in the literature, including DeepSeek-Prover-V1.5 [26], InternLM2.5-StepProver [24], and HunyuanProver [11].
Fixed tactic generation budget. We first evaluate performance under a fixed tactic generation budget, defined as $K\times W\times N$, where $K$ is the total number of passes, $W$ is the expansion width, and $N$ is the maximum number of proof state expansions per pass. This budget represents the total number of LLM inference calls required during proof search. During BFS proof search, BFS-Prover uses a sampling temperature of $1.1$, an expansion width of $2$, and a length normalization factor of $\alpha=0.5$ defined in (1) to maximize the variance of explored proof paths across different search attempts. As shown in Table 1, BFS-Prover achieves state-of-the-art performance under this setting without requiring either a critic model or MCTS-based search. We note that the incorporation of a critic model effectively doubles the number of inference calls since each state expansion requires both policy and value computations.
Accumulative results. In order to maximize potential proof discovery, we conduct BFS searches under multiple hyperparameter configurations and aggregate the successful proofs. Specifically, we vary the BFS length normalization factor $\alpha$ defined in (1) over $\{0.0,0.5,1.0\}$ to execute different exploration and exploitation trade-offs, while other parameter settings remain unchanged. This diverse search strategy allows BFS-Prover to accumulatively achieve a score of 72.95% on the MiniF2F test, setting a new state-of-the-art result in the formal theorem proving literature.
Prover System Critic Search Tactic Budget MiniF2F-test
@@
DeepSeek-Prover-V1.5 [26] No MCTS $32\times 16\times 400$ 63.5%
@@
InternLM2.5-StepProver [24] Yes BFS $256\times 32\times 600$ 65.9%
HunyuanProver [11] Yes BFS $600\times 8\times 400$ 68.4%
BFS-Prover (this work) No BFS $2048\times 2\times 600$ 70.83% $\pm$ 0.89%
BFS-Prover (this work) No BFS accumulative 72.95%
@@
Figure 4: Evolution of proof accuracies with increasing number of proof search passes on the MiniF2F test. We compare two training methods: (blue) only using SFT on all state-tactic pairs accumulated in the expert iteration pipeline of BFS-Prover; and (red), on top of (blue), perform another round of DPO refinement using negative tactic examples described in Section 2.3. It can be seen that DPO achieves consistent improvement in terms of both sample efficiency and overall accuracy.
3.4.2 Scaling Law of BFS and Advantage of Negative Signals from DPO
Finally, we investigate the search-time scaling law of BFS-Prover by examining how its performance on the MiniF2F test benchmark improves with increasing proof search passes and evaluate the advantage of applying DPO to learn from negative signals in enhancing the system. We perform a total of pass@4096 and evaluate intermediate performance at pass@64, pass@128, pass@256, pass@1024, and pass@2048, where the confidence band for each intermediate pass count is computed by sampling multiple pass@64 runs. The experimental results are presented in Fig. 4, where $x$-axis is in the log scale and the shaded region represents min-max range, i.e., the confidence band. In the following, we make several observations about key characteristics of BFS scaling in formal theorem proving. Here, SFT refers to supervised fine-tuning performed on all state-tactic pairs accumulated through the expert iteration pipeline, while SFT+DPO indicates an additional round of DPO refinement applied on top of the SFT model using on-policy negative examples from Lean4 compiler feedback as described in Section 2.3. Both SFT and SFT+DPO employ the following BFS parameter configuration: a sampling temperature of $1.1$, an expansion width of $2$, and a length normalization factor of $\alpha=0.5$.
  • Both SFT and SFT+DPO training methods exhibit logarithmic scaling as the number of proof search passes increases. Specifically, SFT improves from 64.58% to 70.38% and SFT+DPO from 64.98% to 70.83% when scaling passes from 64 to 2048, showing consistent but diminishing returns with doubled computational budget.
  • The SFT+DPO training method consistently outperforms the SFT baseline, demonstrating the effectiveness of incorporating negative feedback from Lean4 compiler errors. This refinement enables the model to better distinguish between successful and unsuccessful proof strategies, leading to improved proof search efficiency and higher success rates.
  • Examining the min-max ranges, both methods exhibit similar variance in performance (around 3-4% range). This indicates that while DPO improves overall success rates, it maintains comparable stability in proof search to the SFT baseline.
4 Conclusion and Discussion